OpenStudy (anonymous):

A stadium has 45000 seats. If baseball fans have bought 5/8 of the seats for the next game, how many seats are available?

OpenStudy (gorv):

bought=(45000)*5/8=5625*5=18125 remain= 45000-18125

OpenStudy (anonymous):

Thanks it worked